# On the robustness of modeling grounded word learning through a child’s egocentric input

### v1

@@ Appendix D Tables
Model CVCL Vong et al. (2024) CVCL CVCL (T) CVCL (T+LM) Architecture Vision encoder ResNeXt-50 ViT-B/16 ViT-B/16 ViT-B/16 Embedding size 512 512 512 512 Language encoder Embedding Embedding Transformer Decoder Transformer Decoder Number of attention heads - - 8 8 Number of layers - - 2 2 Positional encoding - Learned, absolute Learned, absolute Learned, absolute Max sequence length 25 48 48 48 Frame sampling method Sample single frame Sample single frame Sample single frame Sample single frame Pre-trained vision encoder True True True True Fine-tune vision encoder False False False False Image data augmentation True True True True Loss Configuration Loss Contrastive Contrastive Contrastive Contrastive and Language Modeling Fixed temperature True True True True Initial temperature 0.07 0.07 0.07 0.07 $\lambda_{\text{Contrastive}}$ 0.3 Training Configuration Train batch size 8 16 64 64 Validation batch size 8 8 8 8 CLIP-filtered Validation False True True True Optimizer AdamW AdamW AdamW AdamW Learning rate 1e-4 1e-4 1e-5 1e-5 LR Scheduler ReduceLROnPlateau Linear Warmup with Cosine Annealing Linear Warmup with Cosine Annealing Linear Warmup with Cosine Annealing Warmup steps - 5000 5000 5000 Weight decay 0.1 0.1 0.1 0.1 Vision dropout - 0.1 0.1 0.1 Vision layer norm - True True True Language dropout - 0.1 0.1 0.1 Num epochs 400 100 100 100 Early Stopping Validation loss Validation loss Validation loss Validation loss

### v2

@@ Appendix E Tables
Model CVCL Vong et al. (2024) CVCL CVCL (T) CVCL (T+LM) Architecture Vision encoder ResNeXt-50 ViT-B/16 ViT-B/16 ViT-B/16 Embedding size 512 512 512 512 Language encoder Embedding Embedding Transformer Decoder Transformer Decoder Number of attention heads - - 8 8 Number of layers - - 2 2 Positional encoding - Learned, absolute Learned, absolute Learned, absolute Max sequence length 25 48 48 48 Frame sampling method Sample single frame Sample single frame Sample single frame Sample single frame Pre-trained vision encoder True True True True Fine-tune vision encoder False False False False Image data augmentation True True True True Loss Configuration Loss Contrastive Contrastive Contrastive Contrastive and Language Modeling Fixed temperature True True True True Initial temperature 0.07 0.07 0.07 0.07 $\lambda_{\text{Contrastive}}$ - - - 0.3 Training Configuration Train batch size 8 16 64 64 Validation batch size 8 8 8 8 CLIP-filtered Validation False True True True Optimizer AdamW AdamW AdamW AdamW Learning rate 1e-4 1e-4 1e-5 1e-5 LR Scheduler ReduceLROnPlateau Linear Warmup with Cosine Annealing Linear Warmup with Cosine Annealing Linear Warmup with Cosine Annealing Warmup steps - 5000 5000 5000 Weight decay 0.1 0.1 0.1 0.1 Vision dropout - 0.1 0.1 0.1 Vision layer norm - True True True Language dropout - 0.1 0.1 0.1 Num epochs 400 100 100 100 Early Stopping Validation loss Validation loss Validation loss Validation loss
@@
Evaluation Category Labeled-S-V2 Labeled-A Labeled-Y
Frames Videos Frames Videos Frames Videos
Arm 91 42 89 20 67 19
Ball 92 23 36 7 87 11
Banana 71 9 52 4 75 5
Bed 53 18 30 9 43 5
Bedroom 74 13 17 6 49 4
Blanket 60 10 26 5 37 11
Book 83 16 89 9 40 8
Boots 32 4 15 4 69 11
Bottle 86 19 69 9 81 15
Bowl 64 10 51 6 87 13
Box 90 16 75 9 44 5
Bread 34 12 12 1 48 6
Bucket 51 9 13 6 10 4
Cereal 15 4 44 3 77 8
Cow 51 5 14 6 24 1
Cup 65 10 79 8 95 15
Door 74 26 33 12 60 11
Finger 83 46 71 19 57 18
Foot 95 31 79 20 78 18
Hair 30 16 16 6 36 6
Horse 36 4 30 3 12 5
Jeans 71 22 56 11 31 8
Kitchen 95 29 30 8 97 22
Knee 32 11 11 6 10 8
Leg 87 16 83 11 77 9
Monkey 60 6 59 3 15 2
Napkin 29 7 18 3 26 5
Oven 35 7 11 2 37 8
Paper 40 12 19 4 44 5
Pen 10 1 45 4 22 2
Pillow 17 1 42 6 48 5
Porch 92 25 56 7 85 3
Shirt 68 13 21 2 14 9
Shoe 94 15 70 8 75 8
Shoulder 66 21 66 9 21 5
Sock 65 13 52 8 70 15
Sun 54 14 29 10 20 3
Table 83 21 22 8 93 18
Towel 18 9 12 3 10 5
Toy 92 27 91 12 98 20
Window 81 23 87 10 66 11
Table 4: Number of unique exemplars and unique videos per evaluation category in Labeled-S-V2, Labeled-A and Labeled-Y Evaluation Datasets